OpenStudy (anonymous):

Solve 11.32^(x – 1) = 15.7

hero (hero):

\[11.32^{(x – 1)} = 15.7\] log both sides: \[\log11.32^{(x – 1)} = \log15.7\] \[(x - 1)\log(11.32) = \log(15.7)\] \[x - 1 = \frac{\log(15.7)}{\log(11.32)}\] Finish the rest.